Exercise 14-02 a-c

Oriole Corporation was organized on January 1, 2019. During its first year, the corporation issued 1,900 shares of $50 par value preferred stock and 101,000 shares of $10 par value common stock. At December 31, the company declared the following cash dividends: 2019, $5,550; 2020, $14,500; and 2021, $29,000.

Partially correct answer. Your answer is partially correct. Try again.
Show the allocation of dividends to each class of stock, assuming the preferred stock dividend is 7% and noncumulative.

2019

2020

2021

Allocation to preferred stock $

Entry field with correct answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

Allocation to common stock $

Entry field with correct answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

SHOW LIST OF ACCOUNTS

LINK TO TEXT

Partially correct answer. Your answer is partially correct. Try again.
Show the allocation of dividends to each class of stock, assuming the preferred stock dividend is 9% and cumulative.

2019

2020

2021

Allocation to preferred stock $

Entry field with correct answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

Allocation to common stock $

Entry field with incorrect answer now contains modified data

$

Entry field with incorrect answer now contains modified data

$

Entry field with incorrect answer

SHOW LIST OF ACCOUNTS

LINK TO TEXT

Correct answer. Your answer is correct.
Journalize the declaration of the cash dividend at December 31, 2021, under part (b). (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

Dec. 31

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Click if you would like to Show Work for this question:

Open Show Work

SHOW LIST OF ACCOUNTS

SHOW ANSWER

LINK TO TEXT

Question Attempts: 2 of 5 used

SAVE FOR LATER

SUBMIT ANSWER
Answer #1

Show the allocation of dividends to each class of stock, assuming the preferred stock dividend is 7% and noncumulative.

Preferred dividend = 1900*50*7% = 6650

2019

2020

2021

Allocation to preferred stock $5550

Entry field with correct answer

6650 6650
Allocation to common stock $0

Entry field with correct answer

$14500-6650 = 7850

Entry field with incorrect answer

$29000-6650 = 22350

Entry field with incorrect answer

Partially correct answer. Your answer is partially correct. Try again.

Show the allocation of dividends to each class of stock, assuming the preferred stock dividend is 9% and cumulative

Preferred dividend = 1900*50*9% = 8550.

2019

2020

2021

Allocation to preferred stock $5550

Entry field with correct answer

$8550*2-5550 = 11550

Entry field with incorrect answer

$8550

Entry field with incorrect answer

Allocation to common stock $0

Entry field with incorrect answer now contains modified data

$14500-11550 = 2950

Entry field with incorrect answer now contains modified data

$29000-8550 = 20450

Entry field with incorrect answer

Correct answer. Your answer is correct.
Journalize the declaration of the cash dividend at December 31, 2021, under part (b). (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

Dec. 31

Entry field with correct answerCash dividend

Entry field with correct answer29000

Entry field with correct answer

Entry field with correct answerDividend payable

Entry field with correct answer

Entry field with correct answer29000

Click if you would like to Show Work for this question:

Open Show Work